OpenStudy (anonymous):

together, two designers have to make 31 dresses in a month. Janet has to sew 5 more dresses than her partner chris. how many dresses does janet have to make?

OpenStudy (anonymous):

kk well lets say y=chris and x=jane

OpenStudy (anonymous):

x+x=31 y+5=x

OpenStudy (anonymous):

so y=13 and x=a8

OpenStudy (anonymous):

sorry x+x does not equal 31!

OpenStudy (anonymous):

its x+y=31!

OpenStudy (anonymous):

so how many dresses does janet have to make?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

thank you so much!!